how much ammonia gas can be produced from 200 kg of air?


1
Expert's answer
2022-01-10T14:01:28-0500

Molar mass of dry air "=28.9647g\/mol"


Volume "\%" of Nitrogen in dry air "=78.084\%"


Volume ratio"=" Molar ratio


Moles if Nitrogen in "200kg" of air "=\frac{78.084}{100}×\frac{200000}{28.9647}=5391.7" Moles



"N_{2(g)}+3H_{2(g)}\to2NH_{3(g)}"



1 mole of Nitrogen yields 2 moles of Ammonia


"5391.7" moles yields "5391.7\u00d72=10783.4" Moles of Ammonia


Molar mass of Ammonia "=17.031g\/mol"


Ammonia produced "=\frac{10783.4×17.031}{1000}=183.65kg"